# Deflection angle in the strong deflection limit for static axisymmetric spacetimes: local curvature, matter fields, and quasinormal modes

### v2

@@ Abstract
We investigate the deflection of photons in the strong deflection limit within static, axisymmetric spacetimes possessing reflection symmetry. As the impact parameter approaches its critical value, the deflection angle exhibits a logarithmic divergence. This divergence is characterized by a logarithmic rate and a constant offset, which we express in terms of coordinate-invariant curvature evaluated at the unstable photon circular orbit. The curvature contribution is encoded in the electric part of the Weyl tensor, reflecting tidal effects, and the matter contribution is encoded in the Einstein tensor, capturing the influence of local energy and pressure. We also express these coefficients using Newman–Penrose scalars. By exploiting the relationship between the strong deflection limit and quasinormal modes, we derive a new expression for the quasinormal mode frequency in the eikonal limit in terms of the curvature scalars. Our results provide a unified and coordinate-invariant framework that connects observable lensing features and quasinormal modes to the local geometry and matter distribution near compact objects.
@@ I Introduction
Recent observational advances have significantly enhanced our capacity to explore spacetime geometry in the strong-field regime. In particular, the imaging of black hole shadows has vividly illustrated how light propagates near compact objects [1, 2]. At the same time, gravitational lensing has long provided a theoretical basis for understanding the motion of photons through curved spacetimes (see, e.g., Ref. [3] for a review).
@@
When light rays pass extremely close to a compact object, the deflection angle becomes large. In this regime, the strong deflection limit (SDL) provides a useful approximation for analyzing gravitational lensing. The deflection angle diverges logarithmically as the impact parameter approaches its critical value, with a characteristic rate and a constant offset, known as the SDL coefficients. A systematic formulation of this behavior was developed by Bozza [4], and has since been refined and extended in several works [5, 6]. These methods have been widely applied to static, spherically symmetric spacetimes [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], and have further inspired a variety of extensions and applications [25, 26, 27, 28]. Furthermore, this formulation is applicable to the reflection-symmetric plane of axisymmetric spacetimes [29, 30, 31, 32, 33, 34].
@@
Recently, it has been shown that, in static, spherically symmetric spacetimes, the SDL coefficients can be expressed in terms of local geometric and matter field quantities at the photon sphere [35]. This coordinate-independent formulation refines earlier coordinate-based approaches and provides a framework that links local curvature and matter distribution to observable features of strong gravitational lensing. However, whether a similar local and coordinate-independent description remains valid in less symmetric spacetimes has not yet been fully clarified.
@@
Independently of these developments, quasinormal mode (QNM) frequencies have also been found to reflect the properties of unstable photon circular orbits, particularly in the eikonal limit [36, 37]. Interestingly, a connection between the SDL and QNMs was pointed out in previous studies [38, 39], which showed that the imaginary part of the QNM frequency in the eikonal limit is inversely proportional to an SDL coefficient. More significantly, recent work [35] has shown that, in static and spherically symmetric spacetimes, the QNM frequency in the eikonal limit can be expressed in terms of local geometric and matter field quantities evaluated at the photon sphere. This coordinate-invariant formulation provides a physical insight into the QNM frequency as a manifestation of both spacetime curvature and matter distribution near the unstable photon orbit, thereby offering a unified geometric perspective complementary to that of the strong deflection limit.
@@
The relation between QNMs and the SDL is particularly significant in light of recent gravitational wave observations, where the ringdown phase has been detected and matched with theoretical predictions [40, 41, 42]. This underscores the importance of understanding the local geometry near the photon sphere in interpreting strong-field gravitational wave signals.
@@
In this work, we generalize the coordinate-invariant formulation of the deflection angle in the SDL to static and axisymmetric spacetimes with reflection symmetry. We show that the SDL coefficients can be expressed in terms of local geometric and matter field quantities evaluated at the unstable photon circular orbit. Furthermore, we reformulate these coefficients using curvature-based quantities, such as the electric part of the Weyl tensor and Newman–Penrose (NP) scalars and apply the formalism to several explicit spacetimes to illustrate its physical relevance. Using the relation between the SDL and the QNM, we derive a new expression for the QNM frequency in terms of local curvature and matter field quantities.
@@
This paper is organized as follows. In Sec. II, we introduce the general formalism of photon dynamics in static, axisymmetric spacetimes and derive the conditions for the existence of circular photon orbits. In Sec. III, we derive the deflection angle in the SDL by isolating the logarithmically divergent contribution from the photon trajectory integral. In Sec. IV, we recast the SDL expression in a coordinate-invariant form, demonstrating that it depends only on the circumferential radius of the unstable photon circular orbit and local curvature. In Sec. V, we relate this expression to the matter field quantities. In Sec. VI, we further recast the deflection angle in terms of the NP scalars. In Sec. VII, we present several applications of the formalism developed herein. In Sec. VIII, we establish the correspondence between QNM frequencies and SDL coefficients in static, axisymmetric spacetimes. Finally, in Sec. IX, we summarize our findings and discuss their implications for gravitational lensing in the SDL.
@@
II Unstable photon circular orbits in static, axisymmetric spacetimes
We consider a general static, axisymmetric spacetime, which admits Killing fields corresponding to time translations and rotations, adapted to the coordinates $t$ and $\varphi$, respectively. The canonical form of the line element is given by
@@ II Unstable photon circular orbits in static, axisymmetric spacetimes
where $\psi$, $\gamma$, and $W$ are functions of $\rho$ and $\zeta$ [44]. We assume a reflection symmetry (i.e., a $\mathbb{Z}_{2}$ symmetry) with respect to the $\zeta=0$ plane, which requires that $\psi$, $\gamma$, and $W$ are even functions of $\zeta$:
@@
which represents the circumferential radius around the symmetry axis ($\rho=0$). We assume $\rho\in(0,\rho_{\infty})$ and that $R$ diverges as $\rho\to\rho_{\infty}$ for any fixed $\zeta$. Furthermore, on the $\zeta=0$ plane, we assume that the spacetime is asymptotically locally flat; that is, as $\rho\to\rho_{\infty}$, the metric functions approach their Minkowski values: $e^{2\psi}\to 1$, $e^{2\gamma}\to 1$, and $W\to\rho$.
@@
Assuming that $\dot{\varphi}>0$, we divide Eq. (7) by $\dot{\varphi}^{2}$ to yield the radial orbital differential equation:
@@
Next, we consider circular orbits, for which $\dot{\rho}=0$ and $\ddot{\rho}=0$. Consequently, the effective potential $V$ and its derivative $V^{\prime}$, vanish at $\rho=\rho_{\mathrm{m}}$, where $\rho_{\mathrm{m}}$ denotes the coordinate radius of the circular orbit. Setting $V(\rho_{\mathrm{m}})=0$ immediately yields the critical impact parameter $b=b_{\mathrm{c}}$, with
@@
which provides a criterion for the stability of the circular orbits: the orbit is unstable if $V^{\prime\prime}_{\mathrm{m}}<0$, and stable if $V^{\prime\prime}_{\mathrm{m}}>0$. In what follows, we focus on the unstable photon circular orbits (i.e., $V^{\prime\prime}_{\mathrm{m}}<0$), and hence,
$0\displaystyle 2W_{\mathrm{m}}^{2}\psi^{\prime\prime}_{\mathrm{m}}+(W^{\prime}_{\mathrm{m}})^{2}-W_{\mathrm{m}}W^{\prime\prime}_{\mathrm{m}}<$ 0 . (13)
@@ III Deflection angle in the strong deflection limit
Throughout the manuscript, the subscript 00 denotes evaluation at $\rho=\rho_{0}$ and $\zeta=0$ [e.g., $\psi_{0}\equiv\psi(\rho_{0},0)$ and $W_{0}\equiv W(\rho_{0},0)$]. We define the integral
@@
which represents the total angular change experienced by a photon traveling from infinity to $\rho_{0}$ and then back to infinity. Finally, the deflection angle is defined by
@@
$z=1-\frac{R_{0}}{R}.$ (17)
In terms of this variable, the integral $I(\rho_{0})$ becomes
@@
Hence, truncating the expression under the square root in Eq. (18) to second order in $z$ yields $\sqrt{c_{1}z+c_{2}z^{2}}$. Accordingly, we define the corresponding integral as
@@
To express the SDL (i.e., $\rho_{0}\to\rho_{\mathrm{m}}$) in a coordinate-independent manner, we adopt the impact parameter $b$ as a natural measure. This approach directly relates the deviation of $b$ from its critical value $b_{\mathrm{c}}$ to the small difference $(\rho_{0}-\rho_{\mathrm{m}})$. In particular, expanding $b$ around $b_{\mathrm{c}}$ yields
@@
$\bar{b}$ $=\bar{a}\log\left(\frac{2\>\!e^{2(\gamma_{\mathrm{m}}-\psi_{\mathrm{m}})}}{\bar{a}^{2}(R^{\prime}_{\mathrm{m}})^{2}}\right)+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (35)
@@ IV Coordinate-invariant form of the strong deflection coefficients via local curvature
Here, $e_{(0)}^{a}$ represents the four-velocity of static observers. The tetrad components of the Einstein tensor, $G_{(\mu)(\nu)}=G_{ab}e_{(\mu)}^{a}e_{(\nu)}^{b}$, have the following nonzero components:
@@
Using these curvature quantities, the second derivative of the effective potential evaluated at the unstable photon circular orbit [see Eq. (12)] is given by
@@
where the subscript $\mathrm{m}$ of $E_{(\mu)(\nu)}$ and $G_{(\mu)(\nu)}$ indicates evaluation at $\rho=\rho_{\mathrm{m}}$ and $\zeta=0$. For the photon circular orbit to be unstable, we require that $V^{\prime\prime}_{\mathrm{m}}<0$; hence,
@@
This inequality establishes a local, coordinate-invariant relation between the Weyl and Einstein tensors (through their tetrad components) at the unstable photon circular orbit (see also Ref. [46]). In the following section, we reinterpret this inequality in terms of the relationship between the matter field quantities and the tidal forces.
@@
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{12}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!6E_{(3)(3)}^{\mathrm{m}}+G_{(0)(0)}^{\mathrm{m}}-G_{(3)(3)}^{\mathrm{m}}+2\>\!G_{(1)(1)}^{\mathrm{m}}+2\>\!G_{(2)(2)}^{\mathrm{m}}\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (58)
where $R^{\prime}_{\mathrm{m}}\neq 0$. These expressions show that the contributions from $I_{\mathrm{D}}$ to the SDL coefficients depend solely on $R_{\mathrm{m}}$ and the local, coordinate-invariant curvature quantities, $E_{(i)(i)}^{\mathrm{m}}$ and $G_{(\mu)(\mu)}^{\mathrm{m}}$. A comparison with the static, spherically symmetric analysis [35] reveals a key difference: whereas the electric part of the Weyl tensor is absent in that case, it is essential in the present analysis. This underscores the influence of the tidal effects in more general spacetimes.

### v4

# Deflection angle in the strong deflection limit for static and axisymmetric spacetimes: Local curvature, matter fields, and quasinormal modes

@@ Abstract
We investigate the deflection of photons in the strong deflection limit within static and axisymmetric spacetimes possessing reflection symmetry. As the impact parameter approaches its critical value, the deflection angle exhibits a logarithmic divergence. This divergence is characterized by a logarithmic coefficient and a constant offset, which we express in terms of the coordinate-invariant curvature quantities evaluated at the unstable circular photon orbit. The curvature contribution is encoded in the electric part of the Weyl tensor, reflecting tidal effects, and the matter contribution is encoded in the Einstein tensor, capturing the influence of local energy and pressure. We also express these coefficients using the Newman–Penrose scalars. By exploiting the relationship between the strong deflection limit and the quasinormal modes, we derive a new expression for the quasinormal mode frequency in the eikonal limit in terms of the curvature scalars. Our results provide a unified and coordinate-invariant framework that connects observable lensing features and quasinormal modes to the local geometry and matter distribution near compact objects.
@@ I Introduction
Recent observational advances, such as the imaging of black hole shadows by the Event Horizon Telescope [1, 2] and the detection of gravitational waves by LIGO and Virgo [3], have significantly enhanced our capacity to explore spacetime geometry in the strong-field regime. In particular, the imaging of black hole shadows has revealed how light propagates near compact objects, through the observed emission from the surrounding accretion flow. At the same time, gravitational lensing has long provided a theoretical basis for understanding the motion of photons through curved spacetimes (see, e.g., Ref. [4] for a review).
@@
When light rays pass extremely close to a compact object, the deflection angle becomes large. In this regime, the strong deflection limit (SDL) provides a useful approximation for analyzing gravitational lensing. The deflection angle diverges logarithmically as the impact parameter approaches its critical value, with a characteristic rate and a constant offset, known as the SDL coefficients. A systematic description of this behavior was developed by Bozza [5], and has since been refined and extended in several works [6, 7]. These methods have been widely applied to various static and spherically symmetric spacetimes, including the Schwarzschild spacetime [8], the Reissner–Nordström spacetime [9, 10], regular black hole spacetime [11], naked singularity spacetime [12], wormhole spacetimes [17, 16, 13, 14, 15, 18, 19], gravastar spacetimes [20], higher-dimensional black hole spacetimes [21], and black hole spacetimes with quantum corrections [22, 23]. Bozza’s method has also inspired several extensions, such as finite-distance corrections [24, 25], the case of massive particles [26], different source configurations [27], and marginally unstable photon spheres [28]. Furthermore, the same formulation applies to the reflection-symmetric plane of axisymmetric spacetimes, including the Kerr [29, 30], the Kerr–Newman [31, 32], the Majumdar–Papapetrou [33], and the Zipoy–Voorhees spacetimes [34].
@@
Recently, it was shown that, in static and spherically symmetric spacetimes, the SDL coefficients can be expressed in terms of local geometric and matter field quantities at the photon sphere [35]. This coordinate-independent formulation refines earlier coordinate-based approaches and provides a framework that links local curvature and matter distribution to observable features of strong gravitational lensing. However, whether a similar local and coordinate-independent description remains valid in less symmetric spacetimes has not yet been fully clarified.
@@
Independent of these developments, quasinormal mode (QNM) frequencies have also been found to reflect the properties of unstable circular photon orbits in the eikonal limit. Ferrari and Mashhoon [36] first established the correspondence between QNM frequencies and the orbital parameters of the unstable circular photon orbit in this limit. Subsequently, Cardoso et al. [37] generalized this relation to a wide class of static and stationary spacetimes, showing that the imaginary part of the QNM frequency is determined by the Lyapunov exponent of the unstable photon orbit. Stefanov et al. [38] first pointed out a quantitative connection between the SDL and QNMs, showing that the imaginary part of the QNM frequency in the eikonal limit is inversely proportional to the logarithmic coefficient in the SDL. Raffaelli [39] further clarified the geometrical interpretation of this correspondence and discussed its implications for wave propagation near compact objects. More significantly, recent work [35] has shown that, in static and spherically symmetric spacetimes, the QNM frequency in the eikonal limit can be expressed in terms of local geometric and matter field quantities evaluated at the photon sphere. This coordinate-invariant formulation provides a physical insight into the QNM frequency as a manifestation of both spacetime curvature and matter distribution near the unstable photon orbit, thereby offering a unified geometric perspective complementary to that of the strong deflection limit.
@@
The relation between QNMs and the SDL is particularly significant in light of recent gravitational wave observations, where the ringdown phase has been detected and found to be consistent, within the current observational uncertainties, with the theoretical predictions of general relativity [3, 40, 41, 42]. This consistency between the measured and predicted QNM frequencies provides empirical support for the validity of the underlying theoretical framework and motivates further efforts to connect these modes with other observable phenomena in strong gravitational fields. Since the QNM frequencies in the eikonal limit are determined by the properties of unstable photon orbits, this underscores the importance of understanding the local geometry near the photon sphere in interpreting strong-field gravitational wave signals.
@@
In this work, we generalize the coordinate-invariant formulation of the deflection angle in the SDL to static and axisymmetric spacetimes with reflection symmetry. We show that the SDL coefficients can be expressed in terms of local geometric and matter field quantities evaluated at the unstable circular photon orbit. Furthermore, we reformulate these coefficients using curvature-based quantities, such as the electric part of the Weyl tensor and Newman–Penrose (NP) scalars and apply the formalism to several explicit spacetimes to illustrate its physical relevance. Using the relation between the SDL and the QNM, we derive a new expression for the QNM frequency in terms of local curvature and matter field quantities.
@@
This paper is organized as follows. In Sec. II, we introduce the general formalism of photon dynamics in static and axisymmetric spacetimes and derive the conditions for the existence of circular photon orbits. In Sec. III, we derive the deflection angle in the SDL by isolating the logarithmically divergent contribution from the photon trajectory integral. In Sec. IV, we recast the SDL expression in a coordinate-invariant form, demonstrating that it depends only on the circumferential radius of the unstable circular photon orbit and local curvature. In Sec. V, we relate this expression to the matter field quantities. In Sec. VI, we further recast the deflection angle in terms of the NP scalars. In Sec. VII, we present several applications of the formalism developed herein. In Sec. VIII, we establish the correspondence between QNM frequencies and SDL coefficients in static and axisymmetric spacetimes. Finally, in Sec. IX, we summarize our findings and discuss their implications for gravitational lensing in the SDL.
@@
II Unstable circular photon orbits in static and axisymmetric spacetimes
We consider a general static and axisymmetric spacetime, which admits Killing fields corresponding to time translations and rotations, adapted to the time coordinate $t$ and the azimuthal coordinate $\varphi$, respectively. The canonical form of the line element is given by
@@ II Unstable circular photon orbits in static and axisymmetric spacetimes
where $\psi$, $\gamma$, and $W$ are functions of the spatial coordinates $\rho$ and $\zeta$ in the meridional plane, which can be regarded as cylindrical $\rho$-$z$-type coordinates in a particular gauge [44]. We assume a reflection symmetry (i.e., a $\mathbb{Z}_{2}$ symmetry) with respect to the $\zeta=0$ plane, which requires that $\psi$, $\gamma$, and $W$ are even functions of $\zeta$:
@@
which represents the circumferential radius around the symmetry axis ($\rho=0$). We assume $\rho\in(0,\rho_{\infty})$ and that $R$ diverges as $\rho\to\rho_{\infty}$ for any fixed $\zeta$. Furthermore, on the $\zeta=0$ plane, we assume that the spacetime is asymptotically locally flat; that is, as $\rho\to\rho_{\infty}$, the metric functions approach the Minkowski values: $e^{2\psi}\to 1$, $e^{2\gamma}\to 1$, and $W\to\infty$. Note that asymptotic flatness assumed only on the equatorial plane does not guarantee asymptotic flatness off the plane. However, since the following analysis considers the deflection angle of null geodesics confined to the equatorial plane, this assumption is sufficient for our purpose.
@@
Assuming that $\dot{\varphi}\neq 0$, we divide Eq. (7) by $\dot{\varphi}^{2}$ to yield the radial orbital differential equation:
@@
Next, we consider circular orbits, for which $\dot{\rho}=0$ and $\ddot{\rho}=0$. Consequently, the effective potential $V$ and its derivative $V^{\prime}$ vanish at $\rho=\rho_{\mathrm{m}}$, where $\rho_{\mathrm{m}}$ denotes the radius coordinate of the circular orbit. Setting $V(\rho_{\mathrm{m}})=0$ immediately yields the critical impact parameter $b=b_{\mathrm{c}}$ with
@@
which provides a criterion for the stability of the circular orbits: the orbit is unstable if $V^{\prime\prime}_{\mathrm{m}}<0$, and stable if $V^{\prime\prime}_{\mathrm{m}}>0$. In what follows, we focus on the unstable circular photon orbits (i.e., $V^{\prime\prime}_{\mathrm{m}}<0$), and hence,
$2W_{\mathrm{m}}^{2}\psi^{\prime\prime}_{\mathrm{m}}+(W^{\prime}_{\mathrm{m}})^{2}-W_{\mathrm{m}}W^{\prime\prime}_{\mathrm{m}}<0.$ (13)
@@ III Deflection angle in the strong deflection limit
Throughout the manuscript, the subscript 0 denotes evaluation at $\rho=\rho_{0}$ and $\zeta=0$ [e.g., $\psi_{0}\equiv\psi(\rho_{0},0)$ and $W_{0}\equiv W(\rho_{0},0)$]. We define the integral
@@
which represents the total angular change experienced by a photon traveling from infinity to $\rho_{0}$ and then back to infinity. Since the radial coordinate $\rho$ may decrease as it approaches spatial infinity in the chosen coordinate system, the absolute value ensures that the integrand remains positive regardless of whether $\rho$ increases or decreases relative to the turning point $\rho_{0}$. Finally, the deflection angle is defined by
@@
$z=1-\frac{R_{0}}{R},$ (17)
where $R$ is defined in Eq. (3) as a function of $\rho$ and $\zeta$ and is evaluated on the $\zeta=0$ plane. In terms of this variable, the integral $I(\rho_{0})$ becomes
@@
where we have used Eq. (17). Hence, truncating the expression under the square root in Eq. (18) to second order in $z$ yields $\sqrt{c_{1}z+c_{2}z^{2}}$. Accordingly, we define the corresponding integral as
@@
To express the SDL (i.e., $\rho_{0}\to\rho_{\mathrm{m}}$) in a coordinate-independent manner, we adopt the impact parameter $b$ as a natural measure. This approach directly relates the deviation of $b$ from its critical value $b_{\mathrm{c}}$ to the small difference $(\rho_{0}-\rho_{\mathrm{m}})$. In particular, expanding $b$ in Eq. (14) around $\rho_{0}=\rho_{\mathrm{m}}$ and using the circular-orbit condition (11), which eliminates the linear term, we obtain
@@
$\bar{b}$ $=-\bar{a}\log\left[\>\!\frac{\bar{a}^{2}}{2}(R^{\prime}_{\mathrm{m}})^{2}\>\!e^{2(\psi_{\mathrm{m}}-\gamma_{\mathrm{m}})}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (35)
@@ IV Coordinate-invariant form of the strong deflection coefficients via local curvature
Here and hereafter, indices with parentheses, such as $(\mu)$ and $(\nu)$ with $\mu,\nu=0,1,2,3$, denote the components in the orthonormal tetrad basis, whereas unparenthesized indices represent the abstract tensor indices introduced in the Introduction. Here, $e_{(0)}^{a}$ represents the four-velocity of static observers. The tetrad components of the Einstein tensor, $G_{(\mu)(\nu)}=G_{ab}e_{(\mu)}^{a}e_{(\nu)}^{b}$, have the following nonzero components:
@@
Using the curvature quantities defined above, the coordinate derivatives in Eq. (12) can be rewritten in terms of the tetrad components of the curvature tensors given in Eqs. (40)–(52). Applying the $\mathbb{Z}_{2}$ symmetry conditions (54) on the $\zeta=0$ plane and evaluating the result at $\rho=\rho_{\mathrm{m}}$ with the circular-orbit condition (11), we obtain [^1] A brief outline of how Eqs. (39)–(48) are combined to obtain Eq. (55) is presented in Appendix B.
@@
where the subscript $\mathrm{m}$ of $E_{(\mu)(\nu)}$ and $G_{(\mu)(\nu)}$ indicates evaluation at $\rho=\rho_{\mathrm{m}}$ and $\zeta=0$. For the circular photon orbit to be unstable, we require that $V^{\prime\prime}_{\mathrm{m}}<0$; hence,
@@
This inequality establishes a local, coordinate-invariant relation between the Weyl and Einstein tensors (through their tetrad components) at the unstable circular photon orbit (see also Ref. [46]). In the following section, we reinterpret this inequality in terms of the relationship between the matter field quantities and the tidal forces.
@@
$\bar{b}$ $=-\bar{a}\log\left[\>\!\frac{\bar{a}^{2}}{12}R_{\mathrm{m}}^{2}\left(6E_{(3)(3)}^{\mathrm{m}}+G_{(0)(0)}^{\mathrm{m}}-G_{(3)(3)}^{\mathrm{m}}+2\>\!G_{(1)(1)}^{\mathrm{m}}+2\>\!G_{(2)(2)}^{\mathrm{m}}\right)\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (58)
where $R^{\prime}_{\mathrm{m}}\neq 0$. These expressions show that the contributions from $I_{\mathrm{D}}$ to the SDL coefficients depend solely on $R_{\mathrm{m}}$ and the local, coordinate-invariant curvature quantities, $E_{(i)(i)}^{\mathrm{m}}$ and $G_{(\mu)(\mu)}^{\mathrm{m}}$. A comparison with the static and spherically symmetric analysis [35] reveals a key difference: whereas the electric part of the Weyl tensor is absent in that case, it is essential in the present analysis. This underscores the influence of the tidal effects in more general spacetimes.